If the product of two whole numbers is zero, can we say that one or both of them will be zero? Justify through examples.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

If product of two whole numbers is zero, definitely one of them is zero
Example: 0 × 3 = 0 and 15 × 0 = 0
If product of two whole numbers is zero, both of them may be zero
Example: 0 × 0 = 0
Yes, if the product of two whole numbers is zero, then both of them will be zero

Thanks